solve the linear first order equation by using the integrating factor
Solve the linear first order equation (x + 1) = -y +5 using the integrating factor.
Answer #1

x+1 (x+1) dy y = -4+5 s o you + vent = Integrating factor (dx o u(x) = e = m (x+1) = (x+1). Heue solution Aill be e yCx+1) =